OpenStudy (anonymous):

(1/x+2/x+3/x)-(1/a+1/2a+1/3a) When x=2 & a=-2 A. 33/48 B. 39/48 C. 25/12 D. 47/12 When you just say erase x and make it 2 and erase a and make it -2, doesnt really help me at all.

Nnesha (nnesha):

 is your question look like this

Nnesha (nnesha):

@Bri37

OpenStudy (anonymous):

Yes

Nnesha (nnesha):

what answer you got ??

OpenStudy (anonymous):

A

Nnesha (nnesha):

a is not right

Nnesha (nnesha):

can you show me work so i can identify where you mess with that :) so it will be easier for me to explain :)

OpenStudy (anonymous):

(1/2+2/2+3/2)-(1/-2+1/-4+1/6) (6/2)-44/48) (2 44/48)= 2 11/12 2*12=24+11=35/48 I picked A because it was close to my answer

Nnesha (nnesha):

see you forgot to put negative sign with 1/6

Nnesha (nnesha):

 i got this